You have a wire with a cross-sectional area of 0.10 mm2. If there are 4.0 × 1028 atoms per cubic meter in this wire, and if each atom contributes 2 free electrons, what is the drift velocity of the electrons when the current in the wire is 6.0 A?



1
Expert's answer
2021-05-05T14:53:10-0400

to find the drift velocity(due to which electric current is produced in the conductor) of electron

"I" "=" "neAV"d

where "I=" electric current "=6A"

"n=" no. of free electrons per "m"3= charge density = "2 \times 4 \times 10"28  electrons per unit volume

Because each atom contribute 2 free electrons and there are "4 \times 10"28  no. of atoms )

"e=" charge of electron "= 1.6 \times 10"-19

"A=" area of cross section "=0.1 mm"2"= 0.1 \times 10"-6 m2

"V"d "=" drift velocity of electrons

Now put all values in the given above formula

"V"d = "I\/neA = 6\/" "(8 \times 10"28"\times 1.6 \times 10" -19"\times0.1 \times 10"-6")"

"V"d"= 4.7 \times 10"-3 "m\/sec"

so so drift velocity of electron in given conductor is

"= 4.7 \times 10"-3 "m\/sec"